the perimeter of a rectangular swimming pool is 130 yards. Three times the length is equal to ten times the width. what is the length and width of the pool? what is the area?
Answer by checkley77(12844)

3L=10W OR L=10W/3
2L+2W=130
2(10W/3)+2W=130
20W/3+2W=130
(20W+3*2W)/3=130
(20W+6W)/3=130
26W=3*130
26W=390
W=390/26
W=15 ANS. FOR THE WIDTH.
3L=10*15
3L=150
L=150/3
L=50 ANS. FOR THE LENGTH.
PROOF:
2*50+2*15=130
100+30=130
130=130